# GLAD: Global-Local Aware Dynamic Mixture-of-Experts for Multi-Talker ASR

### v1

@@ Abstract
End-to-end multi-talker automatic speech recognition (MTASR) faces significant challenges in accurately transcribing overlapping speech, especially under high-overlap conditions. To address these challenges, we proposed Global-Local Aware Dynamic (GLAD) Mixture-of-Experts, which dynamically fuse speaker-aware global information and fine-grained local features to guide expert selection. This mechanism enables speaker-specific routing by leveraging both global context and local acoustic cues. Experiments on LibriSpeechMix show that GLAD outperforms existing MTASR approaches, particularly in challenging multi-talker scenarios. To our best knowledge, this is the first work to apply Mixture-of-Experts (MoE) to end-to-end MTASR with a global-local fusion strategy. Our code and train dataset can be found at https://github.com/NKU-HLT/GLAD.
@@ 1 Introduction
Motivated by this, we propose a novel Global-Local Aware Dynamic (GLAD) Mixture-of-Experts to enhance MoE for MTASR. Specifically, we introduce a global router that processes raw speech features to capture speaker-aware context. In parallel, local routers model layer-specific information derived from intermediate encoder layers. A dynamic fusion module then adaptively combines global and local routing signals on a per-frame basis, guiding expert selection with both global speaker cues and local acoustic details. Our main contributions are as follows:
@@
(1) First application of MoE to MTASR: To our best knowledge, we are the first to apply MoE architectures to end-to-end MTASR, showing significant improvements over conventional methods.
(2) Novel global-local information fusion: We introduce GLAD, which dynamically combines speaker-aware global context with fine-grained local acoustic features. This enables experts to focus on target speakers while capturing fine-grained speech patterns in multi-talker scenarios.
(3) Interpretability analysis: Our ablation studies investigate the role of global acoustic features and their support for speaker-aware expert routing in MTASR.
@@ 4.3 Case Study
To better understand the behavior of GLAD-SOT, we visualize the cross-attention matrices from the last decoder block on a representative sample (LSM-test-clean-2mix-1840). Figure 2 compares 4 attention heads for SOT-14 and GLAD-SOT. The cross-attention matrices indicate which parts of the input each output token should attend to, revealing the speech segments the model relies on to generate that token. This visualization illustrates the differences in feature modeling between SOT-14 and GLAD-SOT.

### v2

@@ Abstract
End-to-end multi-talker automatic speech recognition (MTASR) faces significant challenges in accurately transcribing overlapping speech, especially under high-overlap conditions. To address these challenges, we proposed Global-Local Aware Dynamic Mixture-of-Experts (GLAD), which dynamically fuse speaker-aware global information and fine-grained local features to guide expert selection. This mechanism enables speaker-specific routing by leveraging both global context and local acoustic cues. Experiments on LibriSpeechMix show that GLAD outperforms existing MTASR approaches, particularly in challenging multi-talker scenarios. To our best knowledge, this is the first work to apply Mixture-of-Experts to end-to-end MTASR with a global-local fusion strategy. Our code and train dataset can be found at https://github.com/NKU-HLT/GLAD.
@@ 1 Introduction
Motivated by this, we propose GLAD (Global-Local Aware Dynamic Mixture-of-Experts), a novel framework designed to improve expert selection in MTASR. Specifically, we introduce a global router that processes raw speech features to capture speaker-aware context. In parallel, local routers model layer-specific information derived from intermediate encoder layers. A dynamic fusion module then adaptively combines global and local routing signals on a per-frame basis, guiding expert selection with both global speaker cues and local acoustic details. We further apply GLAD to SOT paradigm, resulting in GLAD-SOT. Our main contributions are as follows:
@@
(1) To our best knowledge, we are the first to apply MoE architectures to end-to-end MTASR, showing significant improvements over conventional methods.
(2) We introduce GLAD, which dynamically combines speaker-aware global context with fine-grained local acoustic features. This enables experts to focus on target speakers while capturing fine-grained speech patterns in multi-talker scenarios.
(3) Our ablation studies investigate the role of global acoustic features and their support for speaker-aware expert routing in MTASR.
@@ 4.3 Case Study
To better understand the behavior of GLAD-SOT, we visualize the cross-attention matrices from the last decoder block on a representative sample (LSM-test-clean-2mix-1014). Figure 2 compares 4 attention heads for SOT-14 and GLAD-SOT. The cross-attention matrices indicate which parts of the input each output token should attend to, revealing the speech segments the model relies on to generate that token. This visualization illustrates the differences in feature modeling between SOT-14 and GLAD-SOT.